# GenKI: Enhancing Open-Domain Question Answering with Knowledge Integration and Controllable Generation in Large Language Models

### v1

@@ 6.1.2. Can GenKI adapt to different formats of answers?
To demonstrate the adaptability of our model to another format domain, we partition the CMRC dataset, encompassing both long and short-answer formats by their average length. Specifically, we fine-tune our model only on the short-answer segment during step $PostP$, resulting in a noteworthy result of 71.70 on the EM metric. This achievement surpasses all recent baselines, trailing behind only one of the task-specific baselines. Notably, given that these task-specific baselines are exclusively trained intra-domain, our model exhibits exceptional performance, underscoring its remarkable out-of-domain capability, which is shown in Figure 5.
@@
Baseline Type Method EMF1%
Task Specific Baselines ERNIE2.0 69.1 88.6
@@
    $RA$: retriever-augmented baselines, $GA$: generator-augmented baselines (finetuned-LLM baselines)
@@ 6.2.1. To what extent can the acquired knowledge impact the proficiency of the model?
(1) The relationship between retrieval quality and model recall initially exhibited a linear trend across these examples ($R^{2}\textgreater$ > 0.99), but eventually reached a bottleneck state, transitioning into another linear relationship ($R^{2}\textgreater$ > 0.985). This suggests that the model has been extensively integrated with the necessary knowledge. The ultimate bottleneck of the model is correlated with the scale of the model. For instance, the bottleneck for LLaMA-65B occurred at around 90% recall, while the bottleneck for GLM-6B was observed at around 80% recall.
@@
TriviaQA
Metric category K&C K K K&C K K
[…]
@@
$X+K+PostP$ 57.1 64.7 64.9 $X+K+PostP$ 75.9 80.9 82.8
@@
$X+K_{R}+PostP$ 58.0 66.6 67.2 $X+K_{R}+PostP$ 80.5 85.4 86.9
@@
    $PostP$: output after controllable generation, $Rm$: output after reward model, $Ext$: external model selection
@@
Metric category K&C K&C K&C C K&C K K
$X$=GLM-6B BLEU-1 BLEU-2 ROUGE-L Coherence $X$=GLM-6B EMF1% Recall%
@@
$X+K+PostP$ 0.589 0.526 0.455 -4.185 $X+K+PostP$ 77.8 91.0 92.3
@@
$X+K_{R}+PostP$ 0.593 0.529 0.452 -4.198 $X+K_{R}+PostP$ 78.0 91.1 92.2
@@
Figure 5. Comparison of (Left) GenKI, GenKI without knowledge base and best baseline, (Right) GenKI, GenKI in Out-Of-Domain scenario and baselines.
@@ 6.3.1. Can GenKI tackle distribution between knowledge integration or controlled generation?
The success of our model lies in effectively balancing both knowledge integration and controlled generation. To substantiate this claim, we categorize the metrics into two components: knowledge integration metric $K$ and controlled generation metric $C$. As shown in table 5 and table 6 our model succeeds in obtaining all $K\&C$ and $C$ metrics and most $K$ metrics.
We are also concerned about how different structures influence our answers. After undergoing Step $PostP$ in Table 5, the recall of all answers decreases (by up to 2%), while the EM of all answers increases(by up to 2 times). This indicates that the model is more inclined to have learned distribution of instructions which is different from knowledge. This observation underscores the importance of splitting the answer generation process into two steps: knowledge integration and controllable generation. The $Ext$ step plays the same effect on the recall of answers. It performs better when the backbone LLM performs worse. The improvement reaches 18.8% recall in TriviaQA using GLM-6B.

### v2

@@ 6.1.2. Can GenKI adapt to different formats of answers?
To demonstrate the adaptability of our model to another format domain, we partition the CMRC dataset, encompassing both long and short-answer formats by their average length. Specifically, we fine-tune our model only on the short-answer segment during step $\mathrm{PostP}$, resulting in a noteworthy result of 71.70 on the EM metric. This achievement surpasses all recent baselines, trailing behind only one of the task-specific baselines. Notably, given that these task-specific baselines are exclusively trained intra-domain, our model exhibits exceptional performance, underscoring its remarkable out-of-domain capability, which is shown in Figure 5.
@@
Baseline type Method EM% F1%
Task specific
baselines
ERNIE2.0 69.1 88.6
@@
    RARA: retriever-augmented baselines, $GA$: generator-augmented baselines (finetuned-LLM baselines)
@@ 6.2.1. To what extent can the acquired knowledge impact the proficiency of the model?
(1) The relationship between retrieval quality and model recall initially exhibited a linear trend across these examples ($R^{2}\textgreater$ 0.99), but eventually reached a bottleneck state, transitioning into another linear relationship ($R^{2}\textgreater$ 0.985). This suggests that the model has been extensively integrated with the necessary knowledge. The ultimate bottleneck of the model is correlated with the scale of the model. For instance, the bottleneck for LLaMA-65B occurred at around 90% recall, while the bottleneck for GLM-6B was observed at around 80% recall.
@@
XX=GLM-6B EM% F1% Recall% XX=LLaMa-65B EM% F1% Recall%
@@
$X+K+\mathrm{PostP}$ 57.1 64.7 64.9 $X+K+\mathrm{PostP}$ 75.9 80.9 82.8
@@
$X+K_{R}+\mathrm{PostP}$ 58.0 66.6 67.2 $X+K_{R}+\mathrm{PostP}$ 80.5 85.4 86.9
@@
    $\mathrm{PostP}$: output after controllable generation, Rm: output after reward model, Ext: external model selection
@@
XX=GLM-6B BLEU-1 BLEU-2 ROUGE-L Coherence XX=GLM-6B EM% F1% Recall%
@@
$X+K+\mathrm{PostP}$ 0.589 0.526 0.455 -4.185 $X+K+\mathrm{PostP}$ 77.8 91.0 92.3
@@
$X+K_{R}+\mathrm{PostP}$ 0.593 0.529 0.452 -4.198 $X+K_{R}+\mathrm{PostP}$ 78.0 91.1 92.2
@@
Figure 5. Comparison of (Above) GenKI, GenKI without knowledge base and best baseline, (Below) GenKI, GenKI in Out-Of-Domain scenario and baselines.
@@ 6.3.1. Can GenKI tackle distribution between knowledge integration or controlled generation?
[…]
We are also concerned about how different structures influence our answers. After undergoing Step $\mathrm{PostP}$ in table 5, the recall of all answers decreases (by up to 2%), while the EM of all answers increases(by up to 2 times). This indicates that the model is more inclined to have learned distribution of instructions which is different from knowledge. This observation underscores the importance of splitting the answer generation process into two steps: knowledge integration and controllable generation. The $\mathrm{Ext}$ step plays the same effect on the recall of answers. It performs better when the backbone LLM performs worse. The improvement reaches 18.8% recall in TriviaQA using GLM-6B.